undeadknight26 (undeadknight26):

The length of a rectangle is (6x2 + 3x - 2) units, and its width is (x3 - 2x + 5) units. Part A: What is the area of the rectangle? Show your work.

OpenStudy (math_genius123):

area formula: A=L*W

undeadknight26 (undeadknight26):

so: A = (6x^2 + 3x - 2)(x^3 - 2x + 5)

OpenStudy (math_genius123):

yep but solve for doth sides first

OpenStudy (math_genius123):

no first solve for L and W then Multiply them both

undeadknight26 (undeadknight26):

it has a C...but all i need is up 2 b...

OpenStudy (math_genius123):

oh so a and b?

undeadknight26 (undeadknight26):

yup...an i dunno what it is...

OpenStudy (math_genius123):

oh ok just a sec

OpenStudy (math_genius123):

is it 6 multiplied bye 2?

undeadknight26 (undeadknight26):

no its: (6x^2 + 3x - 2)(x^3 - 2x + 5)

OpenStudy (math_genius123):

oh

OpenStudy (math_genius123):

ok

OpenStudy (math_genius123):

dude what grade math is this

undeadknight26 (undeadknight26):

9th...i just really need help man..

OpenStudy (math_genius123):

ok ill try because im like in 7th but i do some collage math

OpenStudy (anonymous):

You have it right undeadknight26: A = (6x^2 + 3x - 2)(x^3 - 2x + 5) A = 6x^5 + 3x^4 - 14x^3 + 24x^2 + 19x -10 Looks like a big derivative

undeadknight26 (undeadknight26):

thanks man :)

OpenStudy (math_genius123):

this might help

undeadknight26 (undeadknight26):

@biire2u does the answer for Part A show that polynomials are closed under an operation?

OpenStudy (math_genius123):

Factor x2y3 + xy I can factor an "x" and a "y" out of each term: x2y3 = xy(x1y2) = xy(xy2) and xy = xy(1). x2y3 + xy = xy( ) = xy(xy2 ) = xy(xy2 + 1) Remember: When "nothing" is left after factoring, a "1" is left behind in the parentheses. Factor 3x3 + 6x2 – 15x. I can factor a "3" and an "x" out of each term: 3x3 = 3x(x2), 6x2 = 3x(2x), and –15x = 3x(–5). Being careful of my signs, I get: 3x3 + 6x2 – 15x = 3x( ) = 3x(x2 ) = 3x(x2 + 2x ) = 3x(x2 + 2x – 5)

OpenStudy (math_genius123):

did u figure it out yet?

undeadknight26 (undeadknight26):

im on part B now...i used birr's answer...

OpenStudy (math_genius123):

oh ok i guess i can go now